# Universal quantum control over bosonic network

### v1

@@ Abstract
Perfect transfer of unknown states across distinct nodes is fundamental in the construction of bosonic quantum networks. We here develop a general theory to control an $N$-node bosonic network governed by the time-dependent Hamiltonian, as the universal quantum control theory for continuous-variable systems. In particular, we can activate nonadiabatic passages superposed of initial and target modes by the commutation condition for the Hamiltonian’s coefficient matrix in the representation of time-independent ancillary modes, which serves as the necessary and sufficient condition to solve the time-dependent Schrödinger equation of the entire network. To exemplify the versatility of our theory on the Heisenberg-picture passages, we perform arbitrary state exchange between two nodes, chiral NOON-state transfer among three bosonic nodes, and chiral Fock-state transfer among three of four bosonic nodes. Our work provides a promising avenue toward universal control of any pair of nodes or modes in bosonic networks as well as the whole network.
@@ I Introduction
Quantum network Kimble (2008); Wehner et al. (2018) exhibits fundamental advantages over its classical counterpart in specific applications, such as quantum key distribution Xu et al. (2020), long-distance quantum computation Duan et al. (2001); Stannigel et al. (2010); Muralidharan et al. (2016); Wallnöfer et al. (2020), distributed quantum computation Lim et al. (2005); Cohen and Mølmer (2018); Cuomo et al. (2020), and quantum metrology Giovannetti et al. (2006); Joo et al. (2011); Giovannetti et al. (2011); Chin et al. (2012); Zhou et al. (2020). A quantum network is typically composed of two or more quantum nodes, constituted of such as atoms Reiserer and Rempe (2015); Liu et al. (2022); Covey et al. (2023), ions Duan and Monroe (2010); Chen et al. (2023), and bosonic modes Briegel et al. (1998); Yung and Bose (2005); Ma et al. (2021); Zhou et al. (2024). State transfer and entanglement distribution in quantum network are mediated by the indirect connection between remote nodes, which can be established via the mutual interaction between neighboring quantum nodes Cirac et al. (1997); Chou et al. (2007); Kimble (2008).
@@
In comparison to the quantum networks constructed by the atomic or ionic nodes, the bosonic network is featured with versatile functions, including (1) the efficient simulation of boson sampling Spring et al. (2013); Broome et al. (2013); Tillmann et al. (2013); Carolan et al. (2014); Spagnolo et al. (2014); Wang et al. (2019); Arrazola et al. (2021), which demonstrates a clear quantum advantage over the classical computer S. and A. (2013), (2) the feasibility of the universal quantum computer based on the Knill-Laflamme-Milburn schemes Knill et al. (2001); Kok et al. (2007), and (3) the fault-tolerant quantum computation with the error-correction codes Chuang et al. (1997); Gottesman et al. (2001); Mirrahimi et al. (2014); Michael et al. (2016). In experiments, bosonic nodes can be set up on multiple platforms, such as cavity quantum electrodynamics (QED) system Kimble (1998); Haroche and Raimond (2006), circuit QED system Vermersch et al. (2017); Blais et al. (2021), hybrid magnonic systems Lachance-Quirion et al. (2019); Xu et al. (2021, 2023); Zheng et al. (2023), synthetic photonic lattices Regensburger et al. (2012); Celi et al. (2014); Lustig et al. (2019); Ozawa and Price (2019), Bose-Einstein condensates Morsch and Oberthaler (2006), and optomechanical systems Heinrich et al. (2011); Stannigel et al. (2011); Aspelmeyer et al. (2014); Zhang et al. (2015); Peterson et al. (2017); Han et al. (2020).
@@
State transfer between bosonic modes has been explored in various protocols of quantum control, particularly those based on the quantum adiabatic theorem Claridge (2009). Under the adiabatic condition, the state transfer between two cavity modes in optomechanical systems Dong et al. (2012); Wang and Clerk (2012); Tian (2012) can be enabled by the mechanical dark mode even under the thermal noise. It mimics the stimulated Raman adiabatic passage in a discrete three-level system Vitanov et al. (2017). Generalized from the three-mode systems Dong et al. (2012); Wang and Clerk (2012); Tian (2012), a dark-mode theorem Huang et al. (2023) recently presents a comprehensive analysis over a bosonic network composed of two coupled subsystems, each of them consisted of multiple individual modes. The evolution along the adiabatically evolved dark mode is however always under a prolonged exposure to environment. It will induce a significant decoherence of the quantum system Jing et al. (2016). Alternatively, a leakage-free path Jing and Wu (2022) enforced by control, such as the general dynamical-decoupling approach Jing et al. (2013, 2015), elevates the condition on achieving adiabatic passage of the system, where the quantum channel is realized through the time-dependent quantum eigenstate. Existing researches suggest that fast and robust state transfer among bosonic modes is usually constrained by the system size, such as the two-mode Zhang et al. (2019); Qi and Jing (2022a); Chen et al. (2018); Li et al. (2021) and three-mode systems Dong et al. (2012); Wang and Clerk (2012); Tian (2012), despite various approaches, such as the transitionless driving Zhang et al. (2019); Qi and Jing (2022a), the inverse engineering Chen et al. (2018), and the pulse optimization Li et al. (2021); Xiang et al. (2023), have inspired the accelerated adiabatic passage in the continuous-variable systems. In general, a universal theoretical framework, which is insensitive to system size, nodes connection (the presence or absence of the dark modes), and target states, is desired for the bosonic networks.
@@
In this paper, we develop a universal theory to control a general $N$-node bosonic network, which extends the universal quantum control (UQC) theory for discrete-variable systems Jin and Jing (2025a, b, c, d, e, f) to that for continuous-variable systems. Universal passages in the Heisenberg picture can be activated by the partial or full commutation condition about the coefficient matrix of the time-dependent network Hamiltonian for the stationary ancillary modes, which is equivalent to solving the time-dependent Schrödinger equation for the whole network. A variety of quantum control over the network of arbitrary size can be performed through the activated passages. Our theory is justified by arbitrary state exchange between two bosonic modes and chiral state transfer across multiple modes.
@@
The rest of this paper is structured as follows. In Sec. II, we introduce a general theoretical framework for solving the time-dependent Schrödinger equation about the bosonic network of arbitrary size. In the representation of stationary ancillary modes, the commutation condition for the coefficient matrix of Hamiltonian gives rise to useful nonadiabatic passages. Section III exemplifies the passage-construction protocol in a paradigmatic two-mode system with mutual state conversion. Section IV presents the chiral NOON-state transfer of three bosonic modes. Section V extends the protocol for a general $N$-node system and demonstrates the chiral Fock-state transfer across four bosonic modes. The entire work is summarized in Sec. VI. Appendix A provides a detailed deduction about the time-dependent and time-independent ancillary-mode transformation for the time-dependent network Hamiltonian.
@@ II General framework
In the framework of UQC theory Jin and Jing (2025a, b, c, d, e), we consider the dynamics of bosonic systems by a set of time-dependent ancillary basis modes $\mu_{k}(t)$’s, $1\leq k\leq N$, which are typically superposed of the bosonic nodes $a_{j}$’s in laboratory. They satisfy the canonical communication relation, i.e., $\mu_{j}(t),\mu_{k}(t)]=\delta_{jk}$.
@@
The adjoint matrix $\mathcal{M}^{\dagger}(t)$ implies the geometric structure of the underlying manifold of $\mu_{k}(t)$’s, admitting a general representation of the form:
@@
$\vec{M}_{1}(t)=(\cos\theta_{1}e^{i\frac{\alpha_{1}}{2}},-\sin\theta_{1}e^{-i\frac{\alpha_{1}}{2}},0,\cdots,0),$ (5)
$\vec{M}_{k}(t)=(\cos\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),-\sin\theta_{k}e^{-i\frac{\alpha_{k}}{2}},0,\cdots,0),$
$\vec{M}_{N-1}(t)=(\cos\theta_{N-1}e^{i\frac{\alpha_{N-1}}{2}}\vec{b}_{N-2}(t),-\sin\theta_{N-1}e^{-i\frac{\alpha_{N-1}}{2}}),$
@@
$\vec{b}_{k}(t)=(\sin\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),\cos\theta_{k}e^{-i\frac{\alpha_{k}}{2}}),$ (6)
where $1\leq k\leq N-1$ and $\vec{b}_{0}(t)\equiv 1$. We can define the bright-mode operators by the inner product $b_{k}(t)=\vec{b}_{k}(t)\cdot\vec{a}_{k+1}^{T}$ with $\vec{a}_{k}^{T}\equiv(a_{1},a_{2},\cdots,a_{k})^{T}$ and $b_{0}(t)=b_{0}(0)\equiv a_{1}$, e.g., $b_{1}(t)=\sin\theta_{1}e^{i\frac{\alpha_{1}}{2}}a_{1}+\cos\theta_{1}e^{-i\frac{\alpha_{1}}{2}}a_{2}$. For the sake of readability, the time-dependence of the parameters $\theta_{k}(t)$ and $\alpha_{k}(t)$ are implicit in Eqs. (5) and (6). They can be either time-dependent or time-independent.
@@
where $H^{\mu}(t)=\mathcal{M}^{\dagger}(t)H^{a}(t)\mathcal{M}(t)$ is the coefficient matrix for $H(t)$ in the basis of the time-dependent ancillary modes $\mu_{k}(t)$’s. To solve the Schrödinger equation with the Hamiltonian in Eq. (7), we have to find a rotation to a representation of time-independent or stationary ancillary modes, by which $\vec{\mu}_{t}\rightarrow\vec{\mu}_{0}$ with $\vec{\mu}_{0}=[\mu_{1}(0),\mu_{2}(0),\cdots,\mu_{N}(0)]$. In general, this rotation can be performed by $V_{N-1}^{\dagger}(t)\mu_{k}(t)V_{N-1}(t)\rightarrow\mu_{k}(0)$ with
@@
with $\delta\alpha_{k}=\alpha_{k}(t)-\alpha_{k}(0)$ and $\delta\theta_{k}=\theta_{k}(t)-\theta_{k}(0)$. The detailed proof can be found in Appendix A.
@@
Main result.— We here prove that the following commutation condition:
@@
where $\Pi^{k}$, $1\leq k\leq N$, is a projection operator or an $N\times N$ coefficient matrix with the only nonzero element at the $k$th row and the $k$th column, i.e., $\Pi^{k}_{jl}=\delta_{jk}\delta_{lk}$, is a necessary and sufficient condition for partially and fully determining the time-evolution operator $U_{\rm rot}(t)$.
@@
with $k$ running from $1$ to $N^{\prime}$, which can be written as Eq. (13) in a more compact form. $N^{\prime}=N$ means that $H_{\rm rot}(t)$ as well as $U_{\rm rot}(t)$ can be fully diagonalized. If both sides of Eq. (15) vanish for arbitrary $t$, then the relevant $\mu_{k}(t)=\mu_{k}(0)$, describing a decoupled dark mode.
Generally, we have
@@
Sufficient condition.— If the coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)]$ for Hamiltonian satisfies the commutation condition in Eq. (13) with $k$ running from $1$ to $N^{\prime}$, then $H_{\rm rot}(t)$ in the relevant subspace takes the diagonal form in Eq. (14). Consequently, the time-evolution operator $U_{\rm rot}(t)$ can be directly given by Eq. (16).
@@
Equation (18) indicates that if the system state presents initially in the mode $\mu_{k}(0)$, then later it will evolve to the mode $\mu_{k}(t)$, with an accumulated global phase $f_{kk}(t)$. In other words, the constraints provided by the commutation condition in Eq. (13) for the time-dependent Hamiltonian $H(t)$ can activate the ancillary mode $\mu_{k}(t)$ as useful nonadiabatic passage. The bosonic modes superposed by the activated passages can be populated through the time evolution under control. In the following sections, the feasibility of our universal control theory can be further verified by the two-mode, three-mode, and $N$-mode bosonic systems.
@@ III State exchange between two bosonic modes
$V_{\alpha_{1}}(t)$ $=e^{-i\frac{\delta\alpha_{1}}{2}(a_{1}^{\dagger}a_{1}-a_{2}^{\dagger}a_{2})},$ (24)
$V_{\theta_{1}}(t)$ $=e^{-\delta\theta_{1}[e^{i\alpha_{1}(0)}a_{2}^{\dagger}a_{1}-e^{-i\alpha_{1}(0)}a_{1}^{\dagger}a_{2}]},$
with $\delta\alpha_{1}=\alpha_{1}(t)-\alpha_{1}(0)$ and $\delta\theta_{1}=\theta_{1}(t)-\theta_{1}(0)$. Substituting Eqs. (20), (21), and (23) to the commutation condition (13), the coupling strength and the detuning can be expressed by the parameters of the ancillary modes in Eq. (21):
@@
The conditions in Eq. (25) share the similar forms as those for a discrete two-dimensional system Jin and Jing (2025a, b, c, d). It is attributed to that the manifold geometry of two bosonic modes is essentially the same as that of a two-level system in determining the gauge field or Berry connection.
According to Eq. (18), the ancillary modes $\mu_{1}(0)$ and $\mu_{2}(0)$ evolve with time as
@@
Equations (21) and (26) show that the ancillary modes $\mu_{1}(t)$ and $\mu_{2}(t)$ can be used to implement the state transfer or exchange between the modes. Both initial and target states can be specified by the proper setting of the boundary conditions, e.g., $\theta_{1}(t)$ and $\alpha_{1}(t)$. For example, arbitrary state of the mode $a_{1}$ can be faithfully transferred to the mode $a_{2}$ via the evolution along the passage $\mu_{1}(t)$ when $t=\tau$, under the conditions of $\theta_{1}(0)=0$ and $\theta_{1}(\tau)=\pi/2$ with $\tau$ the evolution period. During the same period, the initial state of $a_{2}$ is transferred to $a_{1}$ via the passage $\mu_{2}(t)$.
@@
It is straightforward to find that $\Delta=0$ when $\dot{f}=0$, meaning that our passage-construction protocol applied to both degenerate and nondegenerate cases. The latter means that the protocol can directly start from Eq. (19) rather than Eq. (20) in the rotating frame.
@@
In Fig. 1, we demonstrate the performance of our protocol about the state exchange between the modes $a_{1}$ and $a_{2}$ by the fidelity $F=\langle\psi(t)|\rho|\psi(t)\rangle$, where $|\psi(t)\rangle$ is the pure-state solution of the time-dependent Schrödinger equation $id|\psi(t)\rangle/dt=H(t)|\psi(t)\rangle$ with the original Hamiltonian (19). Here $\rho$ can be the initial, intermediate, or target states. In Fig. 1(a), $a_{1}$ is initially prepared in the Fock state $|n=5\rangle$ and $a_{2}$ is prepared in the vacuum state $|0\rangle$. $\rho$ is then chosen such that both modes $a_{1}$ and $a_{2}$ are in Fock states. Consequently, the fidelity can be written as $F_{n_{1},n_{2}}=|\langle n_{1}|\langle n_{2}|\psi(t)\rangle|^{2}$. During the time evolution, the nonadiabatic passage is described by the temporary occupations on the intermediate states: $P_{4,1}=0.410$ when $t=0.30\tau$, $P_{3,2}=0.346$ when $t=0.44\tau$, $P_{2,3}=0.35$ when $t=0.558\tau$, and $P_{1,4}=0.41$ when $t=0.705\tau$. Eventually, the initial Fock state $|5,0\rangle$ completely becomes $|0,5\rangle$ when $t=\tau$. In Fig. 1(b), the initial state is a tensor product of a coherent state and a Fock state $|\alpha=5\rangle_{1}\otimes|n=5\rangle_{2}$. Then $\rho=|\alpha,5\rangle\langle\alpha,5|$ or $\rho=|5,\alpha\rangle\langle 5,\alpha|$. It is found that in the end of the passage, the states of modes $a_{1}$ and $a_{2}$ are perfectly exchanged. And it is interesting to find that in between the beginning and the end of the passage, there exist $n=5$ peaks of fidelity during the time evolution, the same as Fig. 1(a).

### v2

@@ Abstract
Perfect transfer of unknown states across distinct nodes is a basic function in bosonic quantum networks. Here we develop a general theory to construct an $N$-node bosonic network governed by the time-dependent Hamiltonian, as the universal quantum control theory for continuous-variable systems. In particular, we can activate nonadiabatic passages superposed of initial and target modes by the commutation condition about the Hamiltonian’s coefficient matrix and projection operator in the representation of time-independent ancillary modes, which serves as the necessary and sufficient condition to solve the time-dependent Schrödinger equation of the full Hamiltonian. To exemplify the versatility of our theory on the Heisenberg-picture passages, we perform arbitrary state exchange between two nodes, chiral entanglement transfer among three bosonic nodes, and chiral Fock-state transfer among three of four bosonic nodes. Our work provides a promising avenue toward the universal control of any pair of nodes or modes as well as the entire bosonic network.
@@ I Introduction
A quantum network Kimble (2008); Wehner et al. (2018) exhibits fundamental advantages over its classical counterpart in specific applications, such as quantum key distribution Xu et al. (2020), long-distance quantum computation Duan et al. (2001); Stannigel et al. (2010); Muralidharan et al. (2016); Wallnöfer et al. (2020), distributed quantum computation Lim et al. (2005); Cohen and Mølmer (2018); Cuomo et al. (2020), and quantum metrology Giovannetti et al. (2006); Joo et al. (2011); Giovannetti et al. (2011); Chin et al. (2012); Zhou et al. (2020). It is typically composed of two or more quantum nodes, constituted of atoms Reiserer and Rempe (2015); Liu et al. (2022); Covey et al. (2023), ions Duan and Monroe (2010); Chen et al. (2023), and bosonic modes Briegel et al. (1998); Yung and Bose (2005); Ma et al. (2021); Zhou et al. (2024). State transfer and entanglement distribution in a quantum network can be mediated by the indirect connection between remote nodes, which is established via the mutual interactions between neighboring quantum nodes Cirac et al. (1997); Chou et al. (2007); Kimble (2008).
@@
In comparison to the quantum networks constructed by the atomic or ionic nodes, the bosonic network is featured with versatile functions, including (1) the efficient simulation of boson sampling Spring et al. (2013); Broome et al. (2013); Tillmann et al. (2013); Carolan et al. (2014); Spagnolo et al. (2014); Wang et al. (2019); Arrazola et al. (2021), which demonstrates a clear quantum advantage over the classical computer Aaronson and Arkhipov (2013), (2) the feasibility of the universal quantum computer based on the Knill-Laflamme-Milburn schemes Knill et al. (2001); Kok et al. (2007), and (3) the fault-tolerant quantum computation with the error-correction codes Chuang et al. (1997); Gottesman et al. (2001); Mirrahimi et al. (2014); Michael et al. (2016). Bosonic networks can be set up on many physical systems, including a cavity quantum electrodynamics (QED) systems Kimble (1998); Haroche and Raimond (2006), circuit QED systems Vermersch et al. (2017); Blais et al. (2021), synthetic photonic lattices Regensburger et al. (2012); Celi et al. (2014); Lustig et al. (2019); Ozawa and Price (2019), Bose-Einstein condensates Morsch and Oberthaler (2006), optomechanical systems Heinrich et al. (2011); Stannigel et al. (2011); Aspelmeyer et al. (2014); Zhang et al. (2015a); Peterson et al. (2017); Han et al. (2020), circuit quantum acoustodynamics systems von Lüpke et al. (2024), cavity magnomechanical systems Zhang et al. (2016a); Li et al. (2021); Shen et al. (2022, 2025), and cavity magnonic systems Zhang et al. (2014, 2016b); Lachance-Quirion et al. (2019); Xu et al. (2021); Zare Rameshti et al. (2022); Xu et al. (2023); Zheng et al. (2023).
@@
State transfer between bosonic modes or nodes has been explored in various control protocols, particularly those based on the quantum adiabatic theorem Claridge (2009). Under the adiabatic condition, the state transfer between two cavity modes in optomechanical systems Dong et al. (2012); Wang and Clerk (2012); Tian (2012); Fedoseev et al. (2021) can be enabled by the mechanical dark mode even under the thermal noise. It mimics the stimulated Raman adiabatic passage in a discrete three-level system Vitanov et al. (2017). Generalized from the three-mode systems Dong et al. (2012); Wang and Clerk (2012); Tian (2012), a dark-mode theorem Huang et al. (2023) recently presented a comprehensive analysis over a bosonic network composed of two coupled subsystems, each of which consisted of multiple uncoupled modes. The adiabatic evolution along the dark mode is inherently susceptible to the environmental noise due to the long-time exposure. It will induce a significant decoherence of the quantum system Jing et al. (2016). Alternatively, a leakage-free path Jing and Wu (2022) enforced by control, such as the general dynamical-decoupling approach Jing et al. (2013, 2015), elevates the adiabatic condition of the system, where the quantum channel is realized through the time-dependent quantum eigenstate. Existing research suggests that fast and reliable state transfer among bosonic modes is usually constrained by the system size, such as the two-mode Zhang et al. (2019); Qi and Jing (2022a); Chen et al. (2018); Li et al. (2021) and three-mode  Dong et al. (2012); Wang and Clerk (2012); Tian (2012) systems since it is hard to solve the quantum Langevin equation for larger system. Various approaches, such as the transitionless driving Zhang et al. (2019); Qi and Jing (2022a), the inverse engineering Chen et al. (2018), and the pulse optimization Xiang et al. (2023); Lu et al. (2023) seem to inspire the accelerated adiabatic passage in the continuous-variable systems, yet are typically confined to the single-excitation subspace. In general, a universal theoretical framework, which is insensitive to system size, connection geometry (the presence or absence of the dark modes), and target states, is desired for the bosonic networks.
@@
In this paper, we develop a universal theory to control a general $N$-node bosonic network, which extends the universal quantum control (UQC) theory for discrete-variable systems Jin and Jing (2025a, b, c, d, e, f) to that for continuous-variable systems. Universal passages in the Heisenberg picture can be activated by the partial or complete commutation condition about the coefficient matrix of the time-dependent network Hamiltonian in the stationary representation, which is equivalent to solving the time-dependent Schrödinger equation for the entire network. A variety of quantum controls over the network can be performed through the activated passages. Our theory is justified by arbitrary state exchange between two bosonic modes and chiral entanglement state transfer across multiple modes.
@@
The rest of this paper is structured as follows. In Sec. II, we introduce a general theoretical framework for solving the time-dependent Schrödinger equation about the bosonic network of arbitrary size. In the representation of stationary ancillary modes, the commutation condition about the coefficient matrix of the Hamiltonian and the projection operator of matrix bases gives rise to useful nonadiabatic passages. Section III exemplifies the passage-construction protocol in a paradigmatic two-mode system with mutual state conversion. Section IV presents the chiral entanglement transfer of three bosonic modes. Section V describes the protocol for a general $N$-node system and demonstrates the chiral Fock-state transfer across four bosonic modes. Our protocol is found to be robust against the unwanted coupling in the network. Discussion and conclusion can be found in Sec. VI. Appendix A provides a detailed derivation about the ancillary-mode transformation for the time-dependent network Hamiltonian.
@@ II General framework
In the framework of UQC theory Jin and Jing (2025a, b, c, d, e), we consider the dynamics of bosonic systems by a set of time-dependent ancillary basis operators $\mu_{k}(t)$’s, $1\leq k\leq N$, which are typically superposed of the bosonic nodes $a_{j}$’s in laboratory. They satisfy the canonical communication relation, i.e., $\mu_{j}(t),\mu_{k}^{\dagger}(t)]=\delta_{jk}$.
@@
The adjoint matrix $\mathcal{M}^{\dagger}(t)$ implies the geometric structure of the underlying manifold of $\mu_{k}(t)$’s, admitting a general representation of the form
@@
$\vec{M}_{1}(t)$ $=\left(\cos\theta_{1}e^{i\frac{\alpha_{1}}{2}},-\sin\theta_{1}e^{-i\frac{\alpha_{1}}{2}},0,\cdots,0\right),$ (5)
$\vec{M}_{k}(t)$ $=\left[\cos\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),-\sin\theta_{k}e^{-i\frac{\alpha_{k}}{2}},0,\cdots,0\right],$
$\vec{M}_{N-1}(t)$ $=\Big[\cos\theta_{N-1}e^{i\frac{\alpha_{N-1}}{2}}\vec{b}_{N-2}(t),$
$-\sin\theta_{N-1}e^{-i\frac{\alpha_{N-1}}{2}}\Big],$
@@
$\vec{b}_{k}(t)\equiv\left[\sin\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),\cos\theta_{k}e^{-i\frac{\alpha_{k}}{2}}\right],$ (6)
where $1\leq k\leq N-1$ and $\vec{b}_{0}(t)\equiv 1$. We can define the bright-mode operators by the inner product $b_{k}(t)=\vec{b}_{k}(t)\cdot\vec{a}_{k+1}^{T}$ with $\vec{a}_{k}^{T}=(a_{1},a_{2},\cdots,a_{k})^{T}$ and $b_{0}(t)=a_{1}$, e.g., $b_{1}(t)=\sin\theta_{1}e^{i\frac{\alpha_{1}}{2}}a_{1}+\cos\theta_{1}e^{-i\frac{\alpha_{1}}{2}}a_{2}$. For the sake of readability, the time-dependence of the parameters $\theta_{k}(t)$ and $\alpha_{k}(t)$ are implicit in Eqs. (5) and (6). They can be either time-dependent or time-independent.
@@
where $H^{\mu}(t)=\mathcal{M}^{\dagger}(t)H^{a}(t)\mathcal{M}(t)$ is the coefficient matrix for $H(t)$ in terms of the time-dependent ancillary modes $\mu_{k}(t)$’s. To solve the Schrödinger equation with the Hamiltonian in Eq. (7), we have to find a rotation to a representation of time-independent or stationary ancillary modes, by which $\vec{\mu}_{t}\rightarrow\vec{\mu}_{0}$ with $\vec{\mu}_{0}=[\mu_{1}(0),\mu_{2}(0),\cdots,\mu_{N}(0)]$. In general, this rotation can be performed by $V_{N-1}^{\dagger}(t)\mu_{k}(t)V_{N-1}(t)\rightarrow\mu_{k}(0)$ with
@@
with $\delta\alpha_{k}=\alpha_{k}(t)-\alpha_{k}(0)$ and $\delta\theta_{k}=\theta_{k}(t)-\theta_{k}(0)$. The detailed derivation can be found in appendix A.
@@
Main result.— Here we prove that the commutation condition:
@@
is a necessary and sufficient condition for partially and fully determining the time-evolution operator $U_{\rm rot}(t)$. Here $\Pi^{k}$, $1\leq k\leq N$, is a projection operator or an $N\times N$ coefficient matrix with the only nonzero element at the $k$th row and the $k$th column, i.e., $\Pi^{k}_{jl}=\delta_{jk}\delta_{lk}$.
@@
with $k$ running from $1$ to $N^{\prime}$, which can be written as Eq. (13) in a more compact form. $N^{\prime}=N$ means that $H_{\rm rot}(t)$ as well as $U_{\rm rot}(t)$ can be fully diagonalized. If the diagonal element in the gauge potential $\mathcal{A}_{kk}$ vanishes for arbitrary $t$, then the relevant $\mu_{k}(t)=\mu_{k}(0)$ describes a decoupled dark mode due to the fact that $\mathcal{A}_{km}\equiv-i[\mu_{k}(t),d\mu_{m}^{\dagger}(t)/dt]$. Otherwise, the time-dependent ancillary operator $\mu_{k}(t)$ determined by $H^{\mu}(t)$ can activate a useful passage. In general, we have
@@
Sufficient condition.— If the coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ for Hamiltonian satisfies the commutation condition in Eq. (13) with $k$ running from $1$ to $N^{\prime}$, then $H_{\rm rot}(t)$ in the relevant subspace takes the diagonal form in Eq. (14). Consequently, the time-evolution operator $U_{\rm rot}(t)$ can be directly obtained as Eq. (16).
@@
Equation (18) indicates that if the system initial state resides in the mode $\mu_{k}(0)$, then later it will evolve to the mode $\mu_{k}(t)$, with an accumulated global phase $f_{kk}(t)$. In other words, the constraints by the commutation condition in Eq. (13) for the time-dependent Hamiltonian $H(t)$ can activate the ancillary mode $\mu_{k}(t)$ as useful nonadiabatic passage. The bosonic modes appear in the activated passages can be populated through the time evolution under control. In what follows, the feasibility of our universal control theory is further verified by the two-mode, three-mode, and $N$-mode bosonic systems.
@@ III State exchange between two bosonic modes
$V_{\alpha_{1}}(t)$ $=e^{-i\frac{\alpha_{1}(t)-\alpha_{1}(0)}{2}(a_{1}^{\dagger}a_{1}-a_{2}^{\dagger}a_{2})},$ (24)
$V_{\theta_{1}}(t)$ $=e^{-[\theta_{1}(t)-\theta_{1}(0)][e^{i\alpha_{1}(0)}a_{2}^{\dagger}a_{1}-e^{-i\alpha_{1}(0)}a_{1}^{\dagger}a_{2}]}.$
Substituting Eqs. (20), (21), and (23) to the commutation condition (13), the coupling strength and the detuning can be expressed by the parameters of the ancillary modes in Eq. (21):
@@
The conditions in Eq. (25) share the similar forms as those for a discrete two-dimensional system Jin and Jing (2025a, b, c, d). It is attributed to the fact that the manifold geometry of two bosonic modes is essentially the same as that of a two-level system in determining the gauge field or Berry connection.
According to Eq. (18), the ancillary modes $\mu_{1}(0)$ and $\mu_{2}(0)$ in the Heisenberg picture evolve with time as
@@
Equations (21) and (26) show that the ancillary modes $\mu_{1}(t)$ and $\mu_{2}(t)$ can be used to implement the state transfer or exchange. Both initial and target states can be specified by the proper setting of the boundary conditions, e.g., $\theta_{1}(t)$ and $\alpha_{1}(t)$. For example, arbitrary state of the mode $a_{1}$ can be faithfully transferred to the mode $a_{2}$ via the evolution along the passage $\mu_{1}(t)$ when $t=\tau$, under the conditions of $\theta_{1}(0)=0$ and $\theta_{1}(\tau)=\pi/2$ with $\tau$ the evolution period. During the same period, the initial state of $a_{2}$ is transferred to $a_{1}$ via the passage $\mu_{2}(t)$.
@@
It is straightforward to find that the detuning $\Delta(t)$ can be controlled by $\dot{f}(t)$. In particular, $\dot{f}(t)=0$ yields a resonant detuning $\Delta(t)=0$, which is the parallel transport condition in discrete-variable systems as a key element of conventional holonomic transformation Sjöqvist et al. (2012). In contrast, the setting $\dot{f}(t)\neq 0$ gives rise to off-resonance $\Delta(t)\neq 0$. Our passage-construction protocol applies to both resonant and off-resonant detuning regimes. In the latter case, the protocol can directly start from Eq. (19) rather than Eq. (20) in the rotating frame.
@@
In Fig. 1, we demonstrate the performance of our protocol about the state exchange between the modes $a_{1}$ and $a_{2}$ by the fidelity $F=\langle\psi(t)|\rho|\psi(t)\rangle$, where $|\psi(t)\rangle$ is the pure-state solution of the time-dependent Schrödinger equation $id|\psi(t)\rangle/dt=H(t)|\psi(t)\rangle$ with the original Hamiltonian (19). Here $\rho$ can be the initial, intermediate, or target states. In Fig. 1(a), $a_{1}$ is initially prepared in the Fock state $|n=5\rangle$ and $a_{2}$ is prepared in the vacuum state $|0\rangle$. $\rho$ is then chosen such that both modes $a_{1}$ and $a_{2}$ are in Fock states. Consequently, the fidelity can be written as $F_{n_{1},n_{2}}=|\langle n_{1}|\langle n_{2}|\psi(t)\rangle|^{2}$. During the time evolution, the nonadiabatic passage is described by the temporary occupations on the intermediate states: $P_{4,1}=0.41$ when $t=0.30\tau$, $P_{3,2}=0.35$ when $t=0.44\tau$, $P_{2,3}=0.35$ when $t=0.56\tau$, and $P_{1,4}=0.41$ when $t=0.71\tau$. The initial Fock state $|5,0\rangle$ completely becomes $|0,5\rangle$ when $t=\tau$. In Fig. 1(b), the initial state is a tensor product of a coherent state and a Fock state $|\alpha=5\rangle_{1}\otimes|n=5\rangle_{2}$. Then $\rho=|\alpha,5\rangle\langle\alpha,5|$ or $\rho=|5,\alpha\rangle\langle 5,\alpha|$. It is found that in the end of the passage, the states of modes $a_{1}$ and $a_{2}$ are perfectly exchanged. And it is interesting to find that in between the beginning and the end of the passage, there exist $m=5$ fidelity peaks during the time evolution, the same as Fig. 1(a).